OpenStudy (anonymous):

Evaluate the expression below when y = -3. 2y + 7 -------- -4y - 10

OpenStudy (anonymous):

Plug in y = -3 2(-3) +7 --------- -4(-3)-10 = (-6+7 )/ (12 -10) = 1/2